A fine foods store sells excellent Mozart Cheese and sells an average of 64 pounds per month with a known standard deviation of 10 pounds. If the manager would like to meet demand 90% of the time, how many pounds should she start the month with? 76.8 74.2 72.6 71.1
Answer #1

P(X < x) = 0.9

Or, P((X - \mu)/\sigma < (x - \mu)/\sigma) = 0.9

Or, P(Z < (x - 64)/10) = 0.9

Or, (x - 64)/10 = 1.28

Or, x = 1.28 * 10 + 64

Or, x = 76.8

Option - A is correct.